# Neutrino yield and neutron shielding calculations for a high-power target installed in an underground setting

### v1

@@ III-B Shielding
Figure 5 shows a proposed shielding configuration. There is an iron cube with walls 60 cm thick surrounding the source, and the gaps between the sleeve and this cube are also filled with iron. This is surrounded by a further 90 cm thick-walled cube of boron loaded concrete. These thicknesses were found to be adequate in a previous study for the KamLAND site[4], and represent safe limits. There is a further concrete plug in the direction away from the beam entry (which points away from the detector), and a further iron block at the beam entry, which points towards the detector, representing the final bending magnet on the beamline. The outgoing neutron fluxes are recorded on a notional sphere of radius 460 cm around this system.
@@
We now address a number of topics related to shielding and radiation levels in the experimental areas, presenting primarily the methodology for addressing these topics. As stated earlier, our calculations relate specifically to the Yemilab site, however they can be tailored to apply to any experimental site by inserting appropriate local conditions.
III-B1 Local Backgrounds
@@
Figure 6 shows the neutron flux resulting from the shielding arrangement in Figure 5, for various neutron energies, as a function of the angle with respect to the proton direction. Units are neutrons per square mm (on the 460 cm notional sphere) per incident proton: to convert to the perhaps more familiar units of they can be multiplied by .111A typical flux in a large power reactor is of order .
@@
At high energies the neutrons peak in the forward direction, away from the detector. There is a comparatively large flux at large angles, heading towards the detector. These are low energy neutrons emerging from the back face, where the beam hits the target, and going backwards down the beam pipe, perhaps scattering a few times on the way. They are not stopped by the iron in the final bending magnet.
III-B2 Activation of components
@@
A radionuclide inventory shows that the isotopes that are the main contributors are , , , (Table I) although other isotopes like , , etc are also present but at much lower rates.
@@
12 h | 1.88e-16 | , | |
270 d | 2.25e-17 | ||
70 d | 6.01e-17 | ||
1900 d | 3.76e-17 |
The activity of the Cu plate from this simulation is shown in Fig. 7. During beam on operation, additional radiation comes from beam loss (neutrons, rays) or from electronic equipment (X-rays from RF amplifiers). However, when full operation is ongoing, no personnel will be allowed in the area. When the beam is off, there will be a small amount of radiation from activated material which, at least from neutrons emanating from the shielding, will be minimal after several days.
III-B3 Activation in the rock
@@
This depends not only on the escaping neutron flux but on the composition of the surrounding rock and also on the legal definition of “low enough”. We have compared a possible installation at Yemilab in Korea[24] with our previous study which assumed installation at KamLAND in Japan[25].
The regulation for the allowed rate at Yemilab is 10 Bq/g, a factor of 100 bigger than at KamLAND (0.1 Bq/g). An assay of Yemilab rock showed that various types of rocks were found, limestone being the most abundant (almost 90% pure ) followed by quartz porphyry, pyrite and skarn minerals (silicate rocks which contain Ca, Fe, Mn, Al and Si minerals). Unlike the Inishi type rock found at KamLAND, the Yemilab environment is a factor of 10 lower in critical elements [low concentrations of Eu ( ppm ), Co ( ppm) and more important Na ()]. The measured rock composition can be found in Table II.
@@
The rock activation was estimated using the same technique as used for KamLAND [4] but with the Yemilab rock composition. The activation level is measured when the experiment is dismantled (about 4-5 years after the beam is turned off) and any rock activation of isotopes with half lives of less than one year or so would have totally decayed away. The long lived isotopes ( 2-3 years) like , and the pair present in the rock are activated by neutrons penetrating the rock. 22Na is produced by fast neutrons via the (n,2n) channel (11 MeV threshold), while the others are produced by thermal neutron activation. Table III shows a comparison between the abundance of the progenitor isotopes at KamLAND and Yemilab.
Isotop | Half life | Parent | KamLAND | Yemilab |
2.6 y | 6% | 0.022% | ||
5.3 y | 30 ppm | 6 ppm | ||
13.5 y | 3 ppm | 0.1 ppm | ||
8.6 y | 3 ppm | 0.1 ppm |
The simulations included a block of rock placed above the shielding with a 100 cm thickness to evaluate the induced activity at various depths in the rock. The total activity was calculated over the entire rock volume, however the highest activity concentration was measured on the central hot spot. Preliminary studies have shown that even for a shielding configuration consisting of 30 cm Fe and 90 cm of boron loaded concrete the highest central activity was less than Bq/g (Figure 8). This is extremely low compared to the Korean legal requirement ( Bq/g) and proves that rock activation at Yemilab is not an issue. This favourable result is due to several factors. First, the sodium concentration in the rock is very low, as is the concentration of Europium. In addition, the cobalt contamination in the rock is 5 times lower at Yemilab.
III-B4 Tritium production
Tritium is produced in the target system in both the sleeve (by interaction of neutrons with Li) and heavy water coolant. It is a radiation hazard because of its long half life (12 years) and the emission of beta particles. The tritium produced in the sleeve will remain trapped inside and will not be an issue, however the tritium produced in the coolant could be a problem because of potential leaks. It was found that the tritium produced in the heavy water with the Yemilab design is almost four times less than for the KamLAND version, and the exact values per incident proton in both materials are shown in Table IV. These differences result from changes in the target and sleeve geometries.
@@ III-B4 Tritium production
III-B5 Neutron Background in the detector
@@
Neutrons produced by proton interactions in the target system can be a background to analyses if they reach the detector. The analyses are likely to require a MeV prompt signal to reduce confusion with neutron capture, but a fast neutron signal may be above this threshold. Neutron background can be reduced by pulse shape identification that discriminates between protons and electrons or positrons. Nevertheless, it is best to minimise the neutron rate into the detector rather than rely on cuts.
Ideally, the shielding to attenuate these neutrons should reduce the rate in the fiducial volume to below the naturally-occurring neutron level, which results from nuclear reactions induced by high-energy muons that reach the environment of the detector. To the best of our understanding, this is “a few neutrons per year (above 3 MeV) in the fiducial volume.” This represents, consulting the numbers shown in Fig. 10, an overall attenuation factor of about 10-24.
@@
At KamLAND, the shielding consisted of a combination of 2 m Fe and B-rich concrete placed around the target system with an extra 2 m length additional shielding of Fe towards the detector to further shield and prevent the neutrons from entering the detector [4]. While at KamLAND, the shielding design was constrained by the limited space of the cavern, this is not the case at Yemilab where the room that will accommodate the target and shielding system is quite large (a semicircular dome 7 m wide, 7 m tall and 22 m long). The energy spectra of the neutrons produced by the target for different angles with respect to the incident proton beam are shown in Figure 9. The backwards-going neutron flux is greatly reduced and this accounts for the target orientation as fast neutrons must not be allowed to reach the detector (note that the beam undergoes two bends so that it hits the target while moving away from the detector).
@@
The energy spectra of the neutrons going through the vacuum pipe and emanating at the surface of the shielding in the backward direction is shown in Figure 10. To estimate the neutron background reaching the detector, these neutrons were tracked through an additional 5 metres thick iron block, since it is the fast (above 3 MeV) neutrons that can contribute to the additional background. Iron was the material of choice due to its high effective cross section for removing high energy neutrons. These neutrons spread out inside the block of iron, and their intensity is effectively attenuated.
@@
The same technique used for KamLAND simulations, event biasing using particle splitting, was applied to evaluate the neutron flux towards the detector. However, the 5 metres of iron was far too great even when using these variance reduction techniques, and therefore this simulation was done in consecutive stages. The 5 metres block was divided into 10 layers, 50 cm thick. The energy spectrum and spatial distribution of neutrons were recorded at the end of each layer, and a new simulation was started using these values as input parameters for the following Fe layer. This has enabled the simulation of the entire thick block of iron. The energy spectra of the neutrons recorded after each iron layer are shown in Figure 11. The 3 MeV energy cut was applied in all these simulations.
@@
After 5 m of Fe the neutron flux is reduced to neutrons/proton/MeV. This value is much lower than was obtained for the KamLAND shielding. The neutron attenuation in iron recorded in these simulations is shown in Figure 12.
III-B6 Photon background in the detector
@@
There will also be a background due to gamma rays, produced mainly by inelastic processes and also, at much lower rates, by neutron capture and subsequent isotopic decay. Photon inelastic processes and radioactive decay contribute even less. Other sources of photon background include beam loss along the transport line due to gas interactions or loss at collimators or the walls of the beam pipe.
@@
This photon background is not relevant for the detection of particles through IBD events, again as these use the “double flash” coincidence of the prompt positron and neutron capture events, but could in principle be a background for BSM searches which use a “single flash” signal from elastic scattering (ES).
@@
Studies show [26] that the ES process has backgrounds from solar neutrinos and from IBD events with one of the ‘flashes’ undetected, with smaller contributions induced by cosmic rays and radiogenics. Mitigating these requires a visible energy threshold of 3 MeV, and a signal size of 7000 events (over 5 years running) is anticipated.
IV Discussion and conclusions
@@
The design of the IsoDAR target and the sleeve has evolved, with increased understanding of the challenges of target design required to cool the 600 kW target in a confined space. As indicated, the engineering details will be presented in a separate publication. The simulations reported in this paper show that these challenges can be met without compromising the physics reach of the experiment, although high isotopic purity will be required for the lithium in the sleeve.
@@
Photon backgrounds in the detector.
The results obtained give us confidence that such experiments can be successfully deployed without affecting the sensitivity of other experiments in the underground laboratory.
V Acknowledgements
@@
We are grateful to our colleagues from the Center for Underground Physics and Yemilab, in the Republic of Korea who have provided us with detail descriptions of the relevant conditions in their laboratory.

### v2

@@ 3.2 Shielding
Figure 7 shows our designed shielding configuration. The roughly-spherical sleeve is enclosed in a solid iron cylinder (aligned with the beam direction) with an outer lateral diameter of 340 cm. The long axis of the cylinder is also 340 cm. This cylinder is surrounded by boron-loaded concrete, poured into a form that is a cube 520 cm on a side. The minimum concrete thickness is 90 cm along the axes of the cube, and capping the ends of the cylinder. These thicknesses were found to be adequate in a previous study for the KamLAND site [Shieldingpaper], and represent very conservative limits for deployment at Yemilab. The simulation geometry includes a further concrete plug on the right of the figure, in the direction away from the beam entry (which points away from the detector) to block fast neutrons transported down the torpedo cooling pipes from reaching the cavern, and a further iron block at the beam entry, on the side pointing towards the detector, which represents the final bending magnet on the beamline. The simulated outgoing neutron fluxes are recorded on a notional sphere of radius 460 cm around this system. Note, this sphere marks the coordinates of a neutron crossing its surface, but does not record the angle at which it crosses.
@@
In addressing the various topics enumerated above related to shielding and radiation levels in the experimental areas, we present primarily the methodology for addressing these topics. For each topic we collect the relevant appropriate data where available (e.g. the composition of the rock) or clearly state the assumptions we are making where it is not (e.g. the nature of the accelerator components). We use the Geant4 simulations to quantify possible problems, and whether the mitigations we have made are adequate. As stated earlier, our calculations relate specifically to the Yemilab site, however they can be tailored to apply to any experimental site by inserting appropriate local conditions.
3.2.1 Local Backgrounds
@@
Figure 8 shows the neutron density per incident proton resulting from the shielding arrangement in Figure 7 for various neutron energies as a function of , the polar angle with respect to the proton direction, recorded on a sphere of radius 460 cm. The units are neutrons per square mm (on the notional sphere) per incident proton: to convert to the perhaps more familiar units of they can be multiplied by .222A typical flux in a large power reactor is of order .
@@
At high energies the neutrons peak in the forward direction, which, because of the bend in the beamline, is away from the detector. There is a comparatively large flux at large angles, heading towards the detector. These are low energy neutrons emerging from the back face, where the beam hits the target, and going backwards down the beam pipe, perhaps scattering a few times on the way. Neutrons are not stopped by the iron in the final bending magnet, because iron is more or less transparent to neutrons below 3 MeV. However, as mentioned earlier, in this study the neutrons below 3 MeV are disregarded anyway, because they do not lead to background events in the detector.
The step around 120∘ is produced by neutrons emerging from the pipe at cm that can then reach the recording sphere. This feature of the plot can be understood by considering a few typical neutron trajectories, as shown in Figure 9. Note the beam direction matches Figure 8 but is the opposite of Figure 7.
3.2.2 Activation of components
@@
A radionuclide inventory shows that the isotopes that are the main contributors are , , , (Table 1) although other isotopes like , , etc are also present but at much lower rates.
@@
| 12 h | 1.88e-16 | , | |
| 270 d | 2.25e-17 | ||
| 70 d | 6.01e-17 | ||
| 1900 d | 3.76e-17 |
The activity of the Cu plate from this simulation is shown in Fig. 10. During beam on operation, additional radiation comes from beam loss (neutrons, rays) or from electronic equipment (X-rays from RF amplifiers). During operation no personnel will be allowed in the area. When the beam is off, there will be a small amount of radiation from activated material which will be minimal after several days.
Our 60 MeV protons are substantially lower in energy than beams at CERN or Fermilab. In these areas, spallation reactions with very high energy neutrons will activate the air with short lived neutron-deficient isotopes of nitrogen or other light elements, including . These can present significant environmental hazards to personnel, but as stated our energies are too low to produce these activities. Nonetheless, in the spirit of overcautiousness, IsoDAR will plan on keeping its environment at a slight negative pressure, so air from the main (clean) laboratory will always flow into the IsoDAR area, from where it will be pumped directly to the surface.
3.2.3 Activation in the rock
@@
This depends not only on the escaping neutron flux but on the composition of the surrounding rock and also on the legal definition of “low enough". We have compared a possible installation at Yemilab in Korea [Yemilab] with our previous study which assumed installation at KamLAND in Japan [Kamland].
The regulation for the allowed rate at Yemilab is 10 Bq/g, a factor of 100 bigger than at KamLAND (0.1 Bq/g). An assay of Yemilab rock showed that various types of rocks were found, limestone being the most abundant (almost 90% pure ) followed by quartz porphyry, pyrite and skarn minerals (silicate rocks which contain Ca, Fe, Mn, Al and Si minerals). Unlike the Inishi type rock found at KamLAND, the Yemilab environment is a factor of 10 lower in critical elements, with low concentrations of Eu ( ppm ), Co ( ppm) and Na (). The measured rock composition can be found in Table 2.
@@
The activation level will be measured when the experiment will be dismantled (about 4-5 years after the beam is turned off) and any rock activation of isotopes with half lives of less than one year or so would have totally decayed away. The rock activation was estimated using the same technique as used for KamLAND [Shieldingpaper] but with the Yemilab rock composition. The long lived isotopes ( 2-3 years) like , and the pair present in the rock are activated by neutrons penetrating the rock. 22Na is produced by fast neutrons via the (n,2n) channel (11 MeV threshold), while the others are produced by thermal neutron activation. Table 3 shows a comparison between the abundance of the progenitor isotopes at KamLAND and Yemilab.
| Isotope | Half life | Parent | KamLAND | Yemilab |
| 2.6 y | 6% | 0.022% | ||
| 5.3 y | 30 ppm | 6 ppm | ||
| 13.5 y | 3 ppm | 0.1 ppm | ||
| 8.6 y | 3 ppm | 0.1 ppm |
The simulations included a block of rock placed above the shielding with a 100 cm thickness to evaluate the induced activity at various depths in the rock. The total activity was calculated over the entire rock volume, however the highest activity concentration was measured on the central hot spot. Preliminary studies have shown that the induced activity averaged over a central 1 region is 0.028 Bq/g. Even for a shielding configuration consisting of 30 cm Fe and 90 cm of boron loaded concrete the highest central activity was less than Bq/g (Figure 11). This is extremely low compared to the Korean legal requirement ( Bq/g) and proves that rock activation at Yemilab is not an issue. This favorable result is due to several factors. First, the sodium concentration in the rock is very low, as is the concentration of europium. In addition, the cobalt contamination in the rock is 5 times lower at Yemilab.
3.2.4 Tritium production
Tritium is produced in the target system in both the sleeve (by interaction of neutrons with Li) and heavy water coolant. It is a radiation hazard because of its long half life (12 years) and the emission of beta particles. The tritium produced in the sleeve will remain trapped inside and will not be an issue, however the tritium produced in the coolant could be a problem because of potential leaks. It was found that the tritium produced in the heavy water with the Yemilab design is almost four times less than for the KamLAND version, and the exact values per incident proton in both materials are shown in Table 4. These differences result from changes in the target and sleeve geometries.
@@ 3.2.4 Tritium production
3.2.5 Neutron Background in the detector
@@
Neutrons produced by proton interactions in the target system can be a background to analyses if they reach the detector. The physics analyses will require a MeV signal to reduce confusion with neutron capture, but a fast neutron signal may be above this threshold. Neutron background can be reduced by pulse shape identification that discriminates between neutron-induced protons and electrons/positrons. Nevertheless, it is best to minimise the intrinsic neutron rate into the detector rather than rely on cuts.
@@
At KamLAND, the shielding consisted of a combination of 2 m Fe and B-rich concrete placed around the target system with an extra 2 m length additional shielding of Fe towards the detector to further shield and prevent the neutrons from entering the detector [Shieldingpaper]. While at KamLAND, the shielding design was constrained by the limited space of the cavern, this is not the case at Yemilab where the room that will accommodate the target and shielding system is quite large (7 m wide, 7 m maximum height at the peak of a semicircular dome with radius 3.5 meters, and 22 m long). The energy spectra of the neutrons produced by the target for different angles with respect to the incident proton beam are shown in Figure 12. The backwards-going neutron flux is greatly reduced and this accounts for the target orientation as fast neutrons must not be allowed to reach the detector (note that the beam undergoes two bends so that it hits the target while moving away from the detector).
@@
The energy spectra of the neutrons going through the vacuum pipe and emanating at the surface of the shielding in the backward direction is shown in Figure 13. To estimate the neutron background reaching the detector, these neutrons were tracked through an additional 5 metres thick iron block (a nominal 1 m for the last bending magnet and then an additional 4 m) , since it is the fast (above 3 MeV) neutrons that can contribute to the additional background. Iron was the material of choice due to its high effective cross section for removing high energy neutrons. These neutrons spread out inside the block of iron, and their intensity is effectively attenuated.
The picking of 5 meters is not that we view this as our thickness requirement, but as a gauge of whether or not the required attenuation can be achieved at all. As we see, the answer is a resounding Yes, allowing then the optimization of the shielding thickness to match the estimated (or measured) natural flux of fast neutrons at the Yemilab site.
Ideally, the shielding to attenuate these neutrons should reduce the rate in the fiducial volume to below the naturally-occurring neutron level, which results from nuclear reactions induced by high-energy muons that reach the environment of the detector and from natural radioactivity in the environment. To the best of our understanding, this is “a few neutrons per year (above 3 MeV) in the fiducial volume.” This represents, consulting the numbers shown in Fig. 13, an overall attenuation factor of about 10-24.
@@
The same technique used for KamLAND simulations, event biasing using particle splitting, was applied to evaluate the neutron flux towards the detector. However, the 5 metres of iron was far too great even when using these variance reduction techniques, and therefore this simulation was done in consecutive stages. The 5 metres block was divided into 10 layers, 50 cm thick. The energy spectrum and spatial distribution of neutrons were recorded at the end of each layer, and a new simulation was started using these values as input parameters for the following Fe layer. For a correct normalisation of the results, after each stage the results were scaled by a factor equal to the ratio between the number of neutrons per proton that were expected to enter the new shielding layer and the number of neutrons generated in the simulation to produce results with good enough statistics. This has enabled the simulation of the entire thick block of iron, and we find that 500 cm reduces the flux of neutrons by a factor . (The use of consecutive stages in the simulation reduces the statistical error on this to .) The energy spectra of the neutrons recorded after each iron layer are shown in Figure 14. (The origin of the step around 20 MeV in many spectra is unclear: it may or may not be related to the switch between measurements and models in the Shielding Physics List.) The 3 MeV energy cut was applied in all these simulations.
@@
5 m of Fe reduces the neutron rate to neutrons/proton. This value is much lower than was obtained for the KamLAND shielding. The neutron attenuation in iron recorded in these simulations is shown in Figure 15. For the single-flash elastic scattering channel we anticipate a signal rate of events over 4 years [Alonso:2021kyu](the double-flash IBD signal size is much larger) or events per proton, so this shielding is more than adequate to render neutron-induced background negligible.
As the attenuation is so large and the result so important, we have done a simple cross check of our Geant4 results against those of the program MCNPX [MCNP]. For this purpose the full geometry was not modelled: 60 MeV protons were generated in a beryllium slab, creating neutrons. The shielding was represented by successive square iron sheets, 300 cm a side and 10 cm thick, encased in a large cylinder of calcium carbonate. The sampling importance of successive 10 cm sheets increased by a factor of 3.0 which compensates for the attenuation and thus maintains the statistical precision (neutrons crossing the boundary are replaced by 3 neutrons with 1/3 of the weight). Results are shown in Fig. 16. The attenuation is roughly (though not exactly) exponential with an attenuation coefficient ,where the error is estimated from subsampling. This attenuation coefficient is similar but not identical to that from Geant4; in this comparison it should be noted that the MCNPX geometrical model is simpler and that the attenuation is not an exact exponential so a difference is to be expected. Also the MCNPX result depends on the weight multiplier used: a factor of 2.0 is understandably not sufficient, factors of 4.0 or 5.0 predict greater attenuation at large thicknesses. This feature of the MCNP code is not understood, but it probably due to particle weights falling below , the smallest normal single-precision floating point number. In summary, the MCNPX simulation agrees broadly with our Geant4 model, and where it disagrees it does so by predicting even greater attenuation.
3.2.6 Photon background in the detector
@@
There will also be a background due to gamma rays, produced mainly by inelastic processes and also, at much lower rates, by neutron capture and subsequent gamma decay of the excited isotope formed. Photon inelastic processes and radioactive decay contribute even less. Other sources of photon background include beam loss along the transport line due to gas interactions or loss at collimators or the walls of the beam pipe.
@@
This photon background is not relevant for the detection of particles through IBD events, as these use the “double flash” coincidence of the prompt positron and neutron capture events, but could in principle be a background for BSM searches which use a “single flash" signal from (e.g.) elastic scattering.
@@
Studies show [Alonso:2021kyu] that the elastic scattering process has backgrounds from solar neutrinos and from IBD events with one of the ‘flashes’ undetected, with smaller contributions induced by cosmic rays and radiogenics. Mitigating these requires a visible energy threshold of 3 MeV, and a signal size of 7000 events (over 5 years running) is anticipated.
4 Discussion and conclusions
@@
The design of the IsoDAR target and the sleeve has evolved, with increased understanding of the challenges of target design required to cool the 600 kW target in a confined space. Engineering details are presented in a separate publication[PDR]. The simulations reported in this paper show that these challenges can be met without compromising the physics reach of the experiment, although high isotopic purity will be required for the lithium in the sleeve.
@@
Photon backgrounds in the detector
The results obtained encourage us to believe that such experiments can be successfully deployed without affecting the sensitivity of other experiments in the underground laboratory.
5 Acknowledgements
We are grateful to our colleagues from the Center for Underground Physics and Yemilab, in the Republic of Korea who have provided us with detailed descriptions of the relevant conditions in their laboratory.